OpenStudy (anonymous):

10% of 15 litres of orange juice plus 5% of orange juice adds up to 8% of orange juice. How many litres of 5% orange juice was added?

OpenStudy (jh3power):

10/100 *15 + (5/100 )x = 8/100 *(x+15)

OpenStudy (jh3power):

so 3/100 x=3/10 x=10

OpenStudy (anonymous):

what? I don't really get it.

OpenStudy (jh3power):

so... x is the amount of 5% orange juice you get that right?

OpenStudy (anonymous):

yep

OpenStudy (jh3power):

so 10% of 15 litre orange juce makes 1.5 right?

OpenStudy (anonymous):

yes

OpenStudy (jh3power):

same with the other things.

OpenStudy (jh3power):

are you still confused?

OpenStudy (anonymous):

what are the other things again? sorry..

OpenStudy (jh3power):

5 percent of x

OpenStudy (jh3power):

and the sum will make 8percent of x+15

OpenStudy (anonymous):

oh! and? is there more?

OpenStudy (jh3power):

no.. so that's the equation.

OpenStudy (anonymous):

oh, right, dumb me, thanks but how did you make: 10/100*15

OpenStudy (jh3power):

10 percent of 15

OpenStudy (anonymous):

oh, right, right! I get it now! thanks a ton!

OpenStudy (jh3power):

good luck

OpenStudy (anonymous):

back at ya